# Predicting the Performance of Graph Convolutional Networks with Spectral Properties of the Graph Laplacian

### v1

@@ Introduction
For any square matrix $A$, it is the case that $Av=\lambda v$ where $v$ is an eigenvector of the matrix and $\lambda$ is the corresponding eigenvalue. Every square matrix has at least one eigenvalue (Mirchev 2016). In this work, we will be concerned with the eigenvalues of graph Laplacian matrices.
@@
Here $I$ is the identity matrix, the values of $w_{i}$ are the learned parameters of the polynomial filter, and $L^{i}$ denotes raising $L$ to the power $i$ or multiplying $i$ copies of $L$ together (Daigavane, Ravindran, and Aggarwal 2021). There are several types of polynomial filters, with the choice being domain or task specific (Gama et al. 2020).
@@
Let $G=(V,E)$ be an undirected graph with degree matrix $D$, adjacency matrix $A$, and Laplacian L=DAL=D-A. For each vertex $i$, let $v_{i}\in\mathbb{R}^{m}$ denote its feature vector, and let the feature matrix be $V=[v_{1}^{\top}\;\cdots\;v_{n}^{\top}]^{\top}\in\mathbb{R}^{n\times m}$. Then,
@@
Table 1 depicts the performance of GCN in Cora CiteSeer and Polblogs datasets. The accuracy of applying GCN is dropped immediately after applying 3 convolution layers for Cora and CiteSeer dataset and for Polblogs dataset it takes 5 layers. We have preprocessed the PolBlogs dataset by removing communities that have only one member to successfully perform the node classification task. After preprocessing, the number of nodes is 1224 and the classes are 11. Since the dataset does not include node features, we assign each node a unique one-hot encoded vector as its feature representation, resulting in an identity matrix of size $N\times N$, where $N$ is the number of nodes in the preprocessed graph.
@@ A.6 Proof of Lemma 2
By definition, L=DAL=D-A, where $D$ is the degree matrix and $A$ is the adjacency matrix. Since $G$ is undirected, each row of $L$ sums to zero. That is, for all $i\in\{1,\dots,n\}$, we have $\sum_{j=1}^{n}L_{ij}=0$. It follows that $L\mathbf{1}=0$, so $\lambda_{1}=0$ is an eigenvalue of $L$, with eigenvector $\mathbf{1}$. Normalizing $\mathbf{1}$ yields the unit eigenvector $u_{1}=\tfrac{1}{\sqrt{n}}\mathbf{1}$. ∎
@@ A.8 Proof of Lemma 4
$\operatorname{Tr}(\bar{V}^{\top}\bar{V})\;=\;\operatorname{Tr}\!\Bigl{(}\bigl{[}\sum_{i=2}^{n}w_{i}\,\bar{u}_{i}^{\top}\bigr{]}\bigl{[}\sum_{j=2}^{n}\bar{u}_{j}\,w_{j}^{\top}\bigr{]}\Bigr{)}\;=\;\\ \sum_{i=2}^{n}\operatorname{Tr}(w_{i}w_{i}^{\top})\;=\;\sum_{i=2}^{n}\|w_{i}\|_{2}^{2}.$
@@ A.9 Proof of Lemma 5
$\operatorname{Tr}(\bar{V}^{\top}L\bar{V})=\operatorname{Tr}\!\Bigl{(}\Bigl{[}\sum_{i=2}^{n}w_{i}\,u_{i}^{\top}\Bigr{]}L\Bigl{[}\sum_{j=2}^{n}u_{j}\,w_{j}^{\!\top}\Bigr{]}\Bigr{)}\\ =\sum_{i=2}^{n}\sum_{j=2}^{n}w_{i}^{\top}\underbrace{u_{i}^{\top}Lu_{j}}_{=\;\lambda_{j}\delta_{ij}}w_{j}=\sum_{i=2}^{n}\lambda_{i}\,w_{i}^{\top}w_{i}\\ =\sum_{i=2}^{n}\lambda_{i}\,\|w_{i}\|_{2}^{2}.$

### v2

@@ Introduction
For any square matrix AA, it is the case that $Av=\lambda v$ where $v$ is an eigenvector of the matrix and $\lambda$ is the corresponding eigenvalue. Every square matrix has at least one eigenvalue (Mirchev 2016). In this work, we will be concerned with the eigenvalues of graph Laplacian matrices.
@@
Here II is the identity matrix, the values of $w_{i}$ are the learned parameters of the polynomial filter, and $L^{i}$ denotes raising $L$ to the power $i$ or multiplying $i$ copies of $L$ together (Daigavane, Ravindran, and Aggarwal 2021). There are several types of polynomial filters, with the choice being domain or task specific (Gama et al. 2020).
@@
Let $G=(V,E)$ be an undirected graph with degree matrix $D$, adjacency matrix AA, and Laplacian L=DAL=D-A. For each vertex $i$, let $v_{i}\in\mathbb{R}^{m}$ denote its feature vector, and let the feature matrix be $V=[v_{1}^{\top}\;\cdots\;v_{n}^{\top}]^{\top}\in\mathbb{R}^{n\times m}$. Then,
@@
Table 1 depicts the performance of GCN in Cora CiteSeer and Polblogs datasets. The accuracy of applying GCN is dropped immediately after applying 3 convolution layers for Cora and CiteSeer dataset and for Polblogs dataset it takes 5 layers. We have preprocessed the PolBlogs dataset by removing communities that have only one member to successfully perform the node classification task. After preprocessing, the number of nodes is 1224 and the classes are 11. Since the dataset does not include node features, we assign each node a unique one-hot encoded vector as its feature representation, resulting in an identity matrix of size $N\times N$, where NN is the number of nodes in the preprocessed graph.
@@ A.6 Proof of Lemma 2
By definition, L=DAL=D-A, where $D$ is the degree matrix and AA is the adjacency matrix. Since $G$ is undirected, each row of $L$ sums to zero. That is, for all $i\in\{1,\dots,n\}$, we have $\sum_{j=1}^{n}L_{ij}=0$. It follows that $L\mathbf{1}=0$, so $\lambda_{1}=0$ is an eigenvalue of $L$, with eigenvector $\mathbf{1}$. Normalizing $\mathbf{1}$ yields the unit eigenvector $u_{1}=\tfrac{1}{\sqrt{n}}\mathbf{1}$. ∎
@@ A.8 Proof of Lemma 4
$\operatorname{Tr}(\bar{V}^{\top}\bar{V})$ $=$ $\operatorname{Tr}\!\Bigl{(}\bigl{[}\sum_{i=2}^{n}w_{i}\,\bar{u}_{i}^{\top}\bigr{]}\bigl{[}\sum_{j=2}^{n}\bar{u}_{j}\,w_{j}^{\top}\bigr{]}\Bigr{)}$
$=$ $\sum_{i=2}^{n}\operatorname{Tr}(w_{i}w_{i}^{\top})$
$=$ $\sum_{i=2}^{n}\|w_{i}\|_{2}^{2}.$
@@ A.9 Proof of Lemma 5
$\operatorname{Tr}(\bar{V}^{\top}L\bar{V})$ $=$ $\operatorname{Tr}\!\Bigl{(}\Bigl{[}\sum_{i=2}^{n}w_{i}\,u_{i}^{\top}\Bigr{]}L\Bigl{[}\sum_{j=2}^{n}u_{j}\,w_{j}^{\!\top}\Bigr{]}\Bigr{)}$
$=$ $\sum_{i=2}^{n}\sum_{j=2}^{n}w_{i}^{\top}\underbrace{u_{i}^{\top}Lu_{j}}_{=\;\lambda_{j}\delta_{ij}}w_{j}$
$=$ $\sum_{i=2}^{n}\lambda_{i}\,w_{i}^{\top}w_{i}$
$=$ $\sum_{i=2}^{n}\lambda_{i}\,\|w_{i}\|_{2}^{2}.$